A roll of gasket material is 16in wide. What length is needed to obtain 21sq ft of the material?
Answer by greenestamps(13198)



Convert the given information into like units:

16 inches is 16/12 = 4/3 feet.

Area = length * width:

21 = x(4/3)
x = (3/4)*21 = 63/4 = 15 3/4

The required length is 15 3/4 feet, or 15 feet 9 inches.